In a ballistics test, a 25 g bullet traveling horizontally at 1200 m/s goes through a 45 cm thick 350 kg stationary target and emerges with a speed of 1000 m/s. The target is free to slide on a smooth horizontal surface.

(a) How long is the bullet in the target (assume constant acceleration)?
s
What average force does it exert on the target?
N (magnitude only)
(b) What is the target's speed just after the bullet emerges?
Answer #1

Mass of the bullet = m1 = 25 g = 0.025 kg

Mas of the block = m2 = 350 kg

Initial speed of the bullet = V0 = 1200 m/s

Speed of the bullet after coming out of the block = V1 = 1000 m/s

Distance the bullet travels through the block = Thickness of the block = D = 45 cm = 0.45 m

Acceleration of the bullet = a

V12 = V02 + 2aD

(1000)2 = (1200)2 + 2a(0.45)

a = -488888.89 m/s2

Time period the bullet is in the target = T

V1 = V0 + aT

1000 = 1200 + (-488888.89)T

T = 4.09 x 10-4 sec

Magnitude of the average force the bullet exerts on the target = F

F = |m1a|

F = |(0.025)(-488888.89)|

F = 1.222 x 104 N

Speed of the block after the bullet emerges = V2

By conservation of linear momentum,

m1V0 = m1V1 + m2V2

(0.025)(1200) = (0.025)(1000) + (350)V2

V2 = 1.43 x 10-2 m/s

a) Time period the bullet is in the target = 4.09 x 10-4 sec

Magnitude of the average force the bullet exerts on the target = 1.222 x 104 N

b) Target's speed just after the bullet emerges = 1.43 x 10-2 m/s